OpenStudy (anonymous):

(tan theta -1) (sec theta -1) =0

OpenStudy (anonymous):

either \[\tan \theta-1=0\] \[\tan \theta=1,\theta=n \pi+\frac{ \pi }{ 4 }\] where n is an integer.

OpenStudy (anonymous):

or \[\sec \theta-1=0,\sec \theta=1,\frac{ 1 }{ \cos \theta }=1,\cos \theta=1,\theta=2 n \pi\] where n is an integer.